# Beyond Local Sharpness: Communication-Efficient Global Sharpness-aware Minimization

### v1

@@ 6.2 Achieving Local-Global Sharpness Consistency
(a) Local model on class sea
w/ FedGloSS
(b) Local model on class snail w/ FedGloSS
(c) Local model on class sea
w/ FedSmoo
(d) Local model on class snail w/ FedSmoo
To assess the effectiveness of FedGloSS in promoting consistency between local and global loss landscapes, Fig. 5 replicates the analysis previously conducted on FedSam (Fig. 2) for direct comparison. The behavior of local models is shown from both local and global perspectives (“Local loss” and “Global loss”, respectively). Section B.1 offers visualizations for the remaining clients. Compared to FedSam, the gap between local and global loss landscapes in FedGloSS is significantly smaller, and both global and local loss surfaces are found in flat and low-loss regions (Figs. 5(a) and 5(b)). This suggests our method effectively promotes convergence toward aligned low-loss flat regions, minimizing the discrepancy between local and global geometries. This results in a global model residing in a flat minimum in the global landscape (Figs. 6(a) and 6(c)). Figs. 5(c) and 5(d) instead compare FedGloSS with the best-performing SOTA FedSmoo, where the position in the global landscape of FedGloSS’ local models is added for reference. While FedSmoo improves consistency between local and global sharpness compared to FedSam, it falls short of FedGloSS in reaching a flatter global minimum.
Table 1 confirms these claims. By combining ADMM for consistency and server-side Sam for global flatness, FedGloSS prioritizes achieving a flatter global region during training, as proven by the lowest global maximum eigenvalue $\lambda_{1,g}$ and larger $\lambda_{1,l}$, across all clients and methods.
@@ 6.3 Benchmarking FedGloSS against SOTA
Table 3: FedGloSS vs. the state of the art on Cifar datasets, distinguished by local optimizer, Sgd (top) and Sam (bottom), in terms of communication cost and accuracy (%). Best results in bold.
CNN ResNet18
Method Comm. Cifar10 Cifar100 Cifar10 Cifar100
Cost $\alpha=0$ $\alpha=0.05$ $\alpha=0$ $\alpha=0.5$ $\alpha=0.05$ $\alpha=0.5$
Client Sgd FedAvg $1\times$ $59.9\scriptscriptstyle{\pm 0.4}$ $65.7\scriptscriptstyle{\pm 1.0}$ $28.6\scriptscriptstyle{\pm 0.7}$ $38.5\scriptscriptstyle{\pm 0.5}$ $72.6\scriptscriptstyle{\pm 0.1}$ $37.4\scriptscriptstyle{\pm 0.2}$
FedProx $1\times$ $59.8\scriptscriptstyle{\pm 0.5}$ $65.6\scriptscriptstyle{\pm 1.0}$ $28.8\scriptscriptstyle{\pm 0.7}$ $38.7\scriptscriptstyle{\pm}0.4$ $72.2\scriptscriptstyle{\pm 0.2}$ $37.6\scriptscriptstyle{\pm 0.1}$
FedDyn $1\times$ $65.5\scriptscriptstyle{\pm 0.3}$ $70.1\scriptscriptstyle{\pm 1.2}$ $70.2\scriptscriptstyle{\pm 0.6}$ $38.8\scriptscriptstyle{\pm 0.6}$
Scaffold $2\times$ $25.1\scriptscriptstyle{\pm 3.7}$ $54.0\scriptscriptstyle{\pm 2.6}$ $30.0\scriptscriptstyle{\pm 1.1}$ $70.8\scriptscriptstyle{\pm 0.6}$ $38.6\scriptscriptstyle{\pm 0.1}$
[…]
Client Sam FedSam $1\times$ $70.2\scriptscriptstyle{\pm 0.9}$ $71.5\scriptscriptstyle{\pm 1.08}$ $28.7\scriptscriptstyle{\pm 0.5}$ $39.6\scriptscriptstyle{\pm 0.5}$ $72.8\scriptscriptstyle{\pm 0.1}$ $38.5\scriptscriptstyle{\pm 0.1}$
FedDyn $1\times$ $79.3\scriptscriptstyle{\pm 3.1}$ $81.5\scriptscriptstyle{\pm 0.6}$ $72.6\scriptscriptstyle{\pm 0.2}$ $39.6\scriptscriptstyle{\pm 0.8}$
FedSpeed $1\times$ $70.9\scriptscriptstyle{\pm 0.4}$ $72.3\scriptscriptstyle{\pm 1.1}$ $28.9\scriptscriptstyle{\pm 0.5}$ $39.7\scriptscriptstyle{\pm 0.5}$ $72.6\scriptscriptstyle{\pm 0.1}$ $38.8\scriptscriptstyle{\pm 0.6}$
FedGamma $2\times$ $58.9\scriptscriptstyle{\pm 1.8}$ $61.9\scriptscriptstyle{\pm 1.8}$ $29.4\scriptscriptstyle{\pm 1.4}$ $72.2\scriptscriptstyle{\pm 0.1}$ $38.8\scriptscriptstyle{\pm 0.3}$
FedSmoo $2\times$ $81.3\scriptscriptstyle{\pm 0.5}$ $82.8\scriptscriptstyle{\pm 0.6}$ $47.8\scriptscriptstyle{\pm 0.5}$ $51.7\scriptscriptstyle{\pm 0.46}$ $75.3\scriptscriptstyle{\pm 0.6}$ $44.8\scriptscriptstyle{\pm 0.5}$
[…]
This section compares FedGloSS with SOTA methods (Section 6.1) on vision tasks in heterogeneous federated settings. Section B.4 discusses results in homogeneous FL.
6.3.1 FedGloSS on Standard Federated Benchmarks
Table 3 presents results on Cifar100 and Cifar10 with varying levels of heterogeneity on the CNN model. Several observations highlight the advantages of FedGloSS. It is straightforward to notice how FedGloSS achieves the best results among both Sgd and Sam-based approaches while maintaining communication efficiency. FedGloSS with local Sam consistently outperforms the best-performing SOTA FedSmoo by $\approx 2.5$ percentage points in accuracy across all dataset configurations with half the communication cost. FedGloSS also reaches the flattest global minima (e.g., $\lambda_{1}^{\textsc{FedGloSS}}=2.03$ vs. $\lambda_{1}^{\textsc{FedSmoo}}=15.37$ on Cifar10 with $\alpha=0$), as shown in Figs. 7 and 6, achieving the best overall performance. FedGloSS with local Sgd overcomes by $\approx 4$ percentage points all Sgd-based approaches. As studied in (Varno et al., 2022), FedDyn suffers from parameter explosion in highly heterogeneous settings, failing to converge on Cifar100. Differently, FedGloSS successfully employs ADMM to align global and local solutions, with the best results under extreme heterogeneity. Similarly, studies showed Scaffold performs poorly in complex heterogeneous environments (Li et al., 2022; Caldarola et al., 2022), resulting in its inability to converge on Cifar100 alongside FedGamma. Lastly, the last two columns of Table 3 further confirm FedGloSS’ effectiveness, consistently outperforming SOTA methods with the more complex ResNet18 architecture, with $\approx 8$ points higher accuracy w.r.t. FedAvg with both Sgd and Sam, and $+5$ w.r.t. FedSmoo, with the flattest solutions (Figs. 6(e) and 6(f)).
@@ 6.3.1 FedGloSS on Standard Federated Benchmarks
6.3.2 FedGloSS on Real-World Large-Scale Datasets
To further highlight FedGloSS’s effectiveness, we evaluate it on large-scale image classification using the challenging Landmarks-User-160k dataset. LABEL:tab:sota_gld compares FedGloSS with local Sam against the best-performing baselines. FedGloSS is among the few methods, alongside FedSam and FedSmoo, outperforming FedAvg. Similarly to the Cifar100 results (Section 6.3.1), both Scaffold and FedGamma fail to converge. Importantly, FedGloSS achieves the best overall performance ($+3.4\%$ w.r.t. FedAvg) with reduced communication overhead.
@@ 6.3.3 ADMM and SAM Interaction in FedGloSS
ADMM-based methods are often prone to parameter explosion in highly heterogeneous FL settings with many clients Varno et al. (2022). This occurs as multiple gradients accumulate in the global dual variable $\sigma$ (Section 5), causing the parameter norms to grow uncontrollably. However, empirical results indicate that this issue is mitigated with Sam (e.g., see FedDyn vs. FedGloSS in Table 3). We hypothesize this is due to Sam ’s nature: by targeting flat minima, it promotes smaller gradient steps and minimal weight updates, resulting in a more stable algorithm. Fig. 8 confirms our hypothesis by showing Sam’s stability effectively lowers parameter norms and the consequent risk of explosion, particularly when Sam is applied directly to the global model, as in FedGloSS.
@@ 6.3.4 Communication Efficiency with FedGloSS
Communication cost is the main bottleneck in FL (Li et al., 2020a), making its optimization a relevant challenge. As already previously highlighted, FedGloSS considers communication efficiency its primacy concern. Defined $B$ the number of bits exchanged by FedAvg in $T$ training rounds, Table 5 studies FedGloSS’s communication cost against the SOTA baselines in terms of rounds necessary to reach FedAvg’s performance and quantity of exchanged bits. The ADMM-based methods are usually faster, with FedGloSS being the fastest with ResNet18 and MobileNetv2. While FedSmoo is faster when using the CNN model, the transmitted bits double due to its increased communication cost, making FedGloSS the most efficient method in all cases. Analyses on left-out settings in Section B.5.
@@ 6.4.1 Communication-efficient Sharpness
This section studies the efficacy of using the pseudo-gradient from the previous round $\tilde{\Delta}_{\operatorname{\boldsymbol{w}}}^{t-1}$ (Eq. 6) as an estimate of the sharpness measure. FedGloSS uses past gradients as a reliable indication on the global loss landscape and, by aligning global and local optimization paths through ADMM, enables consistent trajectories across rounds.
Figure 9: Loss barriers resulting from interpolating NaiveFedGloSS and FedGloSS’ models, which are found in the same basin. Cifar datasets, CNN.
@@
Table 6 compares FedGloSS with its baseline, NaiveFedGloSS (Section 5), which computes the true perturbation $\operatorname{\hat{\boldsymbol{\epsilon}}}^{t}$ at the expense of doubled communication costs. FedGloSS achieves accuracy comparable to NaiveFedGloSS while maintaining communication efficiency, with minimal or negligible gap in performance. This aligns with the observed sharpness of the achieved minima ($\lambda_{1}$). To ensure a fair comparison in communication cost, Table 6 also compares FedGloSS’ final accuracy with NaiveFedGloSS’ performance at 50% training progress, showing that FedGloSS achieves higher accuracy with the same number of exchanges, benefiting from the additional global optimization steps deriving from its communication-efficient strategy. This shows our approximation does not slow convergence. In addition, our strategy in centralized settings lowers the performance w.r.t. Sam, thus reducing our centralized upper bound w.r.t. NaiveFedGloSS. At equal performance, FedGloSS narrows the gap to the upper bound: $-2.4\%$ on Cifar10 with $\alpha=0$ and $-1.9\%$ with $\alpha=0.05$ vs. respectively $-3.2\%$ and $-2.7\%$ of NaiveFedGloSS w.r.t. Sam. In Cifar100 instead, $-7\%$ on $\alpha=0$ and $-4.2\%$ on $\alpha=0.5$ of FedGloSS vs. $-8.1\%$ and $-5.2\%$ of its baseline. Lastly, our sharpness approximation does not steer the optimization path: models trained with FedGloSS and NaiveFedGloSS end up in the same basin (no loss barrier), with similar flatness (Fig. 9). Aiming to reduce the communication bottleneck while achieving superior performance, these results confirm our choice of FedGloSS over NaiveFedGloSS.
@@ 6.4.2 The Role of Global Consistency and Flatness
Table 7 isolates the impact of global consistency and global sharpness minimization in FedGloSS. We recall FedAvg with client-side Sam is FedSam and using ADMM only for aligning local and global convergence points is FedDyn. Both components significantly impact performance, with their combination leading FedGloSS to the best overall results. FedGloSS is not prone to parameter explosion, achieving the best results even where FedDyn fails to converge (). The flatness of FedGloSS’ solutions w.r.t. FedSam in Fig. 6 confirms the efficacy of its strategy.
@@
7 Discussion
This work tackled the challenge of limited generalization in heterogeneous Federated Learning (FL), prioritizing communication efficiency for real-world use. Building on research linking poor generalization to sharp minima in the loss landscape, we showed data heterogeneity worsens discrepancies between local and global loss surfaces, a problem not resolved by methods focusing only on local sharpness.
To address this issue, we introduced Federated Global Server-side Sharpness (FedGloSS), which finds flat minima in the global loss landscape using server-side Sharpness-Aware Minimization (SAM). FedGloSS achieves communication efficiency by approximating SAM’s sharpness through past global pseudo-gradients, distinguishing it from prior approaches. In addition, by not constraining the clients to use SAM as local optimizer, FedGloSS’ required computational cost can be adapted depending on the local available resources.
Importantly, this work revealed SAM prevents ADMM-related parameter explosion by guiding optimization along flat directions, which reflects in reduced model parameters’ norm, enabling stable updates in heterogeneous FL. A promising future direction could involve exploring alternative approaches to ADMM to align local and global solutions, with the goal of avoiding stateful clients.
Extensive evaluations showed FedGloSS outperforms state-of-the-art methods in accuracy, flatness of the solution and communication efficiency, making it a strong candidate for real-world FL applications.

### v2

@@ 6.2 Achieving Local-Global Sharpness Consistency
To assess the effectiveness of FedGloSS in promoting consistency between local and global loss landscapes, Fig. 4 replicates the analysis previously conducted on FedSam (Fig. 2). The behavior of local models is shown from both local (“Local loss”) and global perspectives (“Global loss”). Sec. B.1 offers visualizations for the remaining clients. Compared to FedSam, the gap between local and global loss landscapes in FedGloSS is significantly smaller, and both global and local loss surfaces are found in flat and low-loss regions (Figs. 4(a) and 4(b)). This suggests our method effectively promotes convergence toward aligned low-loss flat regions, minimizing the discrepancy between local and global geometries. This results in a global model residing in a flat minimum in the global landscape (Figs. 6(a) and 6(c)). Figs. 4(c) and 4(d) instead compare FedGloSS with the best-performing SOTA FedSmoo, where the position in the global landscape of FedGloSS’ local models is added for reference. While FedSmoo improves consistency between local and global sharpness compared to FedSam, it falls short of FedGloSS in reaching a flatter global minimum.
Tab. 1 confirms these claims. By combining ADMM for consistency and server-side Sam for global flatness, FedGloSS prioritizes achieving a flatter global region during training, as proven by the lowest global maximum eigenvalue $\lambda_{1,g}$ and larger $\lambda_{1,l}$, across all clients and methods.
@@ 6.3 Benchmarking FedGloSS against SOTA
Table 3: FedGloSS vs. the state of the art on Cifar datasets, distinguished by local optimizer, Sgd (top) and Sam (bottom), in terms of communication cost and accuracy (%). Best results in bold. Model: CNN.
Method Comm. Cifar10 Cifar100
Cost $\alpha=0$ $\alpha=0.05$ $\alpha=0$ $\alpha=0.5$
Client Sgd FedAvg $1\times$ $59.9\scriptscriptstyle{\pm 0.4}$ $65.7\scriptscriptstyle{\pm 1.0}$ $28.6\scriptscriptstyle{\pm 0.7}$ $38.5\scriptscriptstyle{\pm 0.5}$
FedProx $1\times$ $59.8\scriptscriptstyle{\pm 0.5}$ $65.6\scriptscriptstyle{\pm 1.0}$ $28.8\scriptscriptstyle{\pm 0.7}$ $38.7\scriptscriptstyle{\pm}0.4$
FedDyn $1\times$ $65.5\scriptscriptstyle{\pm 0.3}$ $70.1\scriptscriptstyle{\pm 1.2}$
Scaffold $2\times$ $25.1\scriptscriptstyle{\pm 3.7}$ $54.0\scriptscriptstyle{\pm 2.6}$ $30.0\scriptscriptstyle{\pm 1.1}$
[…]
Client Sam FedSam $1\times$ $70.2\scriptscriptstyle{\pm 0.9}$ $71.5\scriptscriptstyle{\pm 1.08}$ $28.7\scriptscriptstyle{\pm 0.5}$ $39.6\scriptscriptstyle{\pm 0.5}$
FedDyn $1\times$ $79.3\scriptscriptstyle{\pm 3.1}$ $81.5\scriptscriptstyle{\pm 0.6}$
FedSpeed $1\times$ $70.9\scriptscriptstyle{\pm 0.4}$ $72.3\scriptscriptstyle{\pm 1.1}$ $28.9\scriptscriptstyle{\pm 0.5}$ $39.7\scriptscriptstyle{\pm 0.5}$
FedGamma $2\times$ $58.9\scriptscriptstyle{\pm 1.8}$ $61.9\scriptscriptstyle{\pm 1.8}$ $29.4\scriptscriptstyle{\pm 1.4}$
FedSmoo $2\times$ $81.3\scriptscriptstyle{\pm 0.5}$ $82.8\scriptscriptstyle{\pm 0.6}$ $47.8\scriptscriptstyle{\pm 0.5}$ $51.7\scriptscriptstyle{\pm 0.46}$
[…]
This section compares FedGloSS with SOTA methods on vision tasks in heterogeneous FL. Sec. B.4 presents results for $\alpha\in\{1,5,10\}$ on Cifar10 and homogeneous settings.
6.3.1 FedGloSS on Standard Federated Benchmarks
Tab. 3 presents results on Cifar100 and Cifar10 with varying levels of heterogeneity on the CNN model. Several observations highlight the advantages of FedGloSS. FedGloSS achieves the best results among both Sgd and Sam-based approaches while maintaining communication efficiency. FedGloSS with local Sam consistently outperforms the best-performing SOTA FedSmoo by $\approx 2.5$ percentage points in accuracy across all configurations with half the communication cost. Our method reaches the flattest global minima (e.g., $\lambda_{1}^{\textsc{FedGloSS}}=2.03$ vs. $\lambda_{1}^{\textsc{FedSmoo}}=15.37$ on Cifar10 $\alpha=0$), as shown in Figs. 7 and 6, achieving the best overall performance. FedGloSS with local Sgd overcomes by $\approx 4$ percentage points all Sgd-based approaches. FedDyn suffers from parameter explosion in highly heterogeneous settings [56], failing to converge on Cifar100. Differently, FedGloSS successfully employs ADMM to align global and local solutions, with the best results under extreme heterogeneity. Studies showed Scaffold performs poorly in complex heterogeneous environments [33, 4], resulting in its inability to converge on Cifar100 alongside FedGamma. Tab. 5 confirms FedGloSS’ effectiveness, consistently outperforming SOTA methods with the more complex ResNet18 architecture, with $\approx 8$ points higher accuracy w.r.t. FedAvg with both Sgd and Sam, and $+5$ w.r.t. FedSmoo, with the flattest solutions (Figs. 6(e) and 6(f)).
Table 4: ResNet18 on Cifar100
$\alpha=0.5$ and Cifar10 $\alpha=0.05$.
Method Comm. cost C100 Acc. C10 Acc.
Client Sgd FedAvg $1\times$ $37.4\scriptscriptstyle{\pm 0.2}$ $72.6\scriptscriptstyle{\pm 0.1}$
FedProx $1\times$ $37.6\scriptscriptstyle{\pm 0.1}$ $72.2\scriptscriptstyle{\pm 0.2}$
FedDyn $1\times$ $38.8\scriptscriptstyle{\pm 0.6}$ $70.2\scriptscriptstyle{\pm 0.6}$
Scaffold $2\times$ $38.6\scriptscriptstyle{\pm 0.1}$ $70.8\scriptscriptstyle{\pm 0.6}$
[…]
Client Sam FedSam $1\times$ $38.5\scriptscriptstyle{\pm 0.1}$ $72.8\scriptscriptstyle{\pm 0.1}$
FedDyn $1\times$ $39.6\scriptscriptstyle{\pm 0.8}$ $72.6\scriptscriptstyle{\pm 0.2}$
FedSpeed $1\times$ $38.7\scriptscriptstyle{\pm 0.6}$ $72.6\scriptscriptstyle{\pm 0.1}$
FedGamma $2\times$ $38.8\scriptscriptstyle{\pm 0.3}$ $72.2\scriptscriptstyle{\pm 0.1}$
FedSmoo $2\times$ $44.8\scriptscriptstyle{\pm 0.5}$ $75.3\scriptscriptstyle{\pm 0.6}$
[…]
@@ 6.3.1 FedGloSS on Standard Federated Benchmarks
6.3.2 FedGloSS on Real-World Large-Scale Datasets
To further highlight FedGloSS’s effectiveness, we evaluate it on large-scale image classification using the challenging Landmarks-User-160k dataset. Tab. 5 compares FedGloSS with local Sam against the best-performing baselines. FedGloSS is among the few methods, alongside FedSam and FedSmoo, outperforming FedAvg. Similarly to the Cifar100 results (Sec. 6.3.1), both Scaffold and FedGamma fail to converge. Importantly, FedGloSS achieves the best overall performance ($+3.4\%$ w.r.t. FedAvg) with reduced communication overhead.
@@ 6.3.3 ADMM and SAM Interaction in FedGloSS
ADMM-based methods are often prone to parameter explosion in highly heterogeneous FL settings with many clients [56]. This occurs as multiple gradients accumulate in the global dual variable $\sigma$ (Sec. 5), causing the parameter norms to grow uncontrollably. However, empirical results indicate this issue is mitigated with Sam (e.g., see FedDyn vs. FedGloSS in Tab. 3). We attribute this to Sam ’s nature: by targeting flat minima, it promotes smaller gradient steps and minimal weight updates, resulting in a more stable algorithm. Fig. 8 confirms our hypothesis by showing Sam’s stability effectively lowers parameter norms and the consequent risk of explosion, particularly when Sam is applied directly to the global model, as in FedGloSS.
@@ 6.3.4 Communication Efficiency with FedGloSS
Communication cost is the main bottleneck in FL [32], making its optimization a relevant challenge. As already previously highlighted, FedGloSS considers communication efficiency its primacy concern. Defined $B$ the number of bits exchanged by FedAvg in $T$ training rounds, Tab. 6 studies FedGloSS’s communication cost against the SOTA baselines in terms of rounds necessary to reach FedAvg’s performance and quantity of exchanged bits. The ADMM-based methods are usually faster, with FedGloSS being the fastest with ResNet18 and MobileNetv2. While FedSmoo is faster when using the CNN model, the transmitted bits double due to its increased communication cost, making FedGloSS the most efficient method in all cases. Analyses on left-out settings in Sec. B.5. Fig. 9 further shows the total communication cost of the FL algorithms (# model exchanges per round ×$\times|\mathcal{C}^{t}|\times\,T$). The doubled cost of FedSmoo is clear, while FedGloSS achieves nearly double FedAvg’s accuracy at the same cost.
@@
Figure 9: Accuracy gain vs. communication cost. CNN, Cifar10 $\alpha=0$.
@@ 6.4.1 Communication-efficient Sharpness
This section evaluates using the pseudo-gradient from the previous round $\tilde{\Delta}_{\operatorname{\boldsymbol{w}}}^{t-1}$ (Eq. 6) to estimate sharpness. FedGloSS leverages past gradients as a reliable indication on the global loss landscape and aligns global and local paths via ADMM for consistent trajectories across rounds.
Tab. 7 compares FedGloSS with its baseline, NaiveFedGloSS (Sec. 5), which computes the exact perturbation $\operatorname{\hat{\boldsymbol{\epsilon}}}^{t}$ at the expense of doubled communication costs. FedGloSS matches NaiveFedGloSS in accuracy while maintaining communication efficiency, with minimal or negligible gap in performance, consistent with similar sharpness ($\lambda_{1}$). To fairly assess communication costs, Tab. 7 also reports FedGloSS’ final accuracy vs. NaiveFedGloSS’ performance at 50% training progress, showing FedGloSS’ higher accuracy given the same number of exchanges. Overall, given $B$ bits transmitted by FedGloSS, the cost of NaiveFedGloSS is $1.96B$ on C10 $\alpha=0$, $2B$ for $\alpha=0.05$, $1.94B$ on C100 $\alpha=0$ and $1.97B$ for $\alpha=0.5$. Thus, FedGloSS reduces communication costs by up to 50%, with no trade-off in model performance. If NaiveFedGloSS were preferred to maintain exact gradients, it would still achieve the best results with costs lower or comparable to FedSmoo. In addition, our strategy in centralized settings lowers the performance w.r.t. Sam, thus reducing our centralized upper bound w.r.t. NaiveFedGloSS. At equal performance, FedGloSS narrows the gap to the upper bound: $-2.4\%$ on Cifar10 with $\alpha=0$ and $-1.9\%$ with $\alpha=0.05$ vs. respectively $-3.2\%$ and $-2.7\%$ of NaiveFedGloSS w.r.t. Sam. In Cifar100 instead, $-7\%$ on $\alpha=0$ and $-4.2\%$ on $\alpha=0.5$ of FedGloSS vs. $-8.1\%$ and $-5.2\%$ of its baseline. Aiming to reduce communication bottlenecks and improve performance, these results validate choosing FedGloSS over NaiveFedGloSS. Further analyses in Sec. B.5.
@@ 6.4.2 The Role of Global Consistency and Flatness
Tab. 8 isolates the impact of global consistency and global sharpness minimization in FedGloSS. We recall FedAvg with client-side Sam is FedSam and using ADMM only for aligning local and global convergence points is FedDyn. Both components significantly impact performance, with their combination leading FedGloSS to the best overall results. FedGloSS is not prone to parameter explosion, achieving the best results even where FedDyn fails to converge (). The flatness of FedGloSS’ solutions w.r.t. FedSam in Fig. 6 confirms the efficacy of its strategy.
@@
7 Conclusion
This work tackled the challenge of limited generalization in heterogeneous Federated Learning (FL), prioritizing communication efficiency for real-world use. Building on research linking poor generalization to sharp minima in the loss landscape, we showed data heterogeneity worsens discrepancies between local and global loss surfaces, a problem not resolved by methods focusing only on local sharpness. To address this, our Federated Global Server-side Sharpness (FedGloSS) finds flat minima in the global loss landscape with server-side Sharpness-Aware Minimization and achieves communication efficiency by approximating SAM’s sharpness through past global pseudo-gradients, distinguishing it from prior approaches. This work revealed SAM prevents ADMM-related parameter explosion by guiding optimization along flat directions, enabling stable updates in heterogeneous FL. Extensive evaluations showed FedGloSS outperforms SOTA methods in accuracy, flatness and communication efficiency.
Acknowledgments.
This study was carried out within the FAIR - Future Artificial Intelligence Research and received funding from the European Union Next-GenerationEU (PIANO NAZIONALE DI RIPRESA E RESILIENZA (PNRR) – MISSIONE 4 COMPONENTE 2, INVESTIMENTO 1.3 – D.D. 1555 11/10/2022, PE00000013). This manuscript reflects only the authors’ views and opinions, neither the European Union nor the European Commission can be considered responsible for them. We acknowledge the CINECA award under the ISCRA initiative for the availability of high-performance computing resources and support.